40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 11 02:25:11 2012, yuta, Update, LSC, PRMI beam spot motion might be from PRM/BS motion POPDCASDC4.pngFeedback4.pngOLPITYAWandPOPDC4.pngOLPITYAWandPOPDCunlocked.png
    Reply  Wed Jul 11 03:53:41 2012, yuta, Update, LSC, BS 3.3 Hz motion on MI REFLDCASDCMIlocked.png
Message ID: 6954     Entry time: Wed Jul 11 02:25:11 2012     Reply to this: 6955
Author: yuta 
Type: Update 
Category: LSC 
Subject: PRMI beam spot motion might be from PRM/BS motion 

My hypothesis from the measurements below, to explain PRMI beam spot motion is;

  Stack motion at 3.3 Hz largely couples to BS and PRM angular motion.
  LSC for PRMI try to compensate this 3.3 Hz motion because they appear in the error signal.
  But since it's not length, failing and even adding more angular motion.

Some plots:
  1. Uncalibrated spectra of POPDC and ASDC when PRMI is locked. This tells you that beam motion seen at POP is 3.3 Hz.

  2. Uncalibrated spectra of feedback signal to BS and PRM. This tells you that LSC is actuating BS and PRM mainly at 3.3 Hz. I think this is because beam spot on PD moves at 3.3 Hz and so faking the error signal.

  3. Below left is uncalibrated spectra of BS, ITMX, ITMY, PRM (and ETMY) angular motion measured using oplevs. I centered oplevs on these optics (except ETMY, which was mis-aligned during PRMI lock). It looks like BS and PRM motion at 3.3 Hz is larger than other optics. Also, there's some coherence between POPDC and BS/PRM motion. We see some coherence with ITMs and even with ETMY, which is completely independent from PRMI. I think this is because 3.3 Hz motion is originated from the ground (stack) motion.

  left:  OLPITYAWandPOPDC4.png          right: OLPITYAWandPOPDCunlocked.png

  4. Above right is the same spectra, but when PRMI is not locked. It looks like there's no big change compared with PRMI locked. When locked, there's some excess for BS and PRM at ~1-3 Hz. I think this is from LSC feedback, which in principle, doesn't affect any angular motion.

Next:
  - Why BS and PRM has large 3.3 Hz peak compared with other optics?
  - Is 3.3 Hz peak effecting MI lock or arm lock?
  - How can we monitor PR2/3 angular motion?

Attachment 1: POPDCASDC4.png  30 kB  | Hide | Hide all | Show all
POPDCASDC4.png
Attachment 2: Feedback4.png  18 kB  | Show | Hide all | Show all
ELOG V3.1.3-